OpenStudy (anonymous):

write the definite integral that represents the area of the surface formed by revolving the graph f(x)=x^(1/2) on the interval [0,4] about the y-axis (Do not evaluate the integral.)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (watchmath):

\[\int_0^2 2\pi y^2\sqrt{1+(2y)^2}\,dy\]

OpenStudy (anonymous):

because it is root x, wouldn't it turn in to x, not y squared?

OpenStudy (watchmath):

\(y=x^{12}\) is equivalent to \(x=y^2\)

OpenStudy (watchmath):

I mean \(y={x^{1/2}}\) is equivalent to \(x=y^2\)

OpenStudy (watchmath):

hold on.... I think I made a mistake

OpenStudy (watchmath):

So \(y=x^{1/2}\) and \(dy/dx=(1/2)x^{-1/2}\). So the surface area is \[\int_0^42\pi x\sqrt{1+(1/4)x^{-1}}\, dx\]